# Power allocation for cell-free MIMO integrated sensing and communication

### v1

@@ IX Conclusions
This paper presents, for the first time, a comprehensive investigation of the power allocation problem in a cell-free (CF) MIMO-ISAC framework, where joint location and velocity estimation is considered as the sensing task. We employ newly derived approximate Cramér-Rao lower bounds (CRLBs) as sensing performance metrics, alongside user receive SINR as the communication metric. To address the inherent nonlinearity and non-convexity of the resulting optimization problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS). Additionally, we introduce a penalty function-based normalized conjugate gradient algorithm with inexact line search (P-NCG-ILS) to handle the power minimization problem in pure sensing scenarios. We analyze the convergence behavior and qualitatively compare the computational complexity of all proposed methods. Simulation results confirm the accuracy of the derived CRLBs and demonstrate the efficacy of the proposed power allocation strategies in enhancing pure sensing and overall ISAC performance. The findings offer valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
Appendix A Proof of Lemma 1
@@
It is obvious that $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when $i>I$ if and only if either ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$. On the other hand, in light of (46), we have
$\Delta L({\bm{\rho}}_{i+1})\approx\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}.$ (65)
@@
Firstly, if ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ for $i>I$, we readily have $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$. Assume there exists $i>I$, where $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ with neither ${\bm{d}}_{i}\rightarrow{\bm{0}}$ nor $v_{i}\rightarrow 0$. We can solve the step size for $\Delta L({\bm{\rho}}_{i+1})=0$ as
$\upsilon_{i}=\dfrac{-2{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}},$ (66)
@@
Appendix B Proof of Theorem 1
@@
$\Longrightarrow\ \upsilon_{i}\leq\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}.$ (67)
Where positive ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ is considered. Assuming the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded, indicating a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus non-zero $\upsilon_{i}$. Furthermore, in light of (48), we have,
$\Delta^{2}L({\bm{\rho}}_{i+1})\leq\epsilon_{L}^{2}\upsilon_{i}^{2}({\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}))^{2}.$ (68)
Since $L({\bm{\rho}})$ is a monotonously decreasing and bounded function regarding ${\bm{\rho}}_{i}$, we have $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$. The deflection factor computed by (45) or the descent direction vector in (53) can guarantee the strict descent ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$. Consequently, sequence $\{L({\bm{\rho}}_{i})\}$ will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$ when it has Lipschitz gradient continuity. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when ${\bm{d}}_{i}\rightarrow{\bm{0}}$. On the other hand, we assume the objective function is not Lipschitz gradient continuous. In this case, if the sequence $\{{\bm{\rho}}_{i}\}$ pass a Lipschitz noncontinuous point ${\bm{\rho}}_{I}$ where ${\bm{d}}_{I}^{\rm T}\nabla^{2}L({\bm{\rho}}_{I}){\bm{d}}_{I}$ is extremely large, to satisfy (67), the step size should meet $\upsilon_{I}\rightarrow 0$. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ when $\upsilon_{i}\rightarrow 0$ for $i\geq I$, i.e., the sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ converge to a Lipschitz gradient noncontinuous point. $\blacksquare$

### v2

@@ VIII Conclusions
This paper presents, for the first time, a comprehensive investigation of the power allocation problem in a CF MIMO-ISAC framework, where joint location and velocity estimation were considered as the sensing task. We employed newly derived approximate CRLBs as sensing performance metrics, alongside user receive SINR as the communication metric. To address the inherent nonlinearity and non-convexity of the resulting optimization problem, we propose the PP-MCG-ILS algorithm. Additionally, the P-NCG-ILS algorithm was further proposed to handle the power minimization problem in pure sensing scenarios. We devoted into the detailed analysis of the convergence behavior and qualitatively compare the computational complexity of all proposed methods. Simulation results confirmed the accuracy of the derived CRLBs and demonstrate the efficacy of the proposed power allocation strategies in enhancing pure sensing and overall ISAC performance. Our findings offer valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
Appendix A Relevant parameters of the derived CRLBs
The waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, are given by
$\small[a_{n,k},\ b_{n,k},\ d_{n,k}]=\rho_{n}\big{[}\underbrace{8\pi^{2}S\bar{f_{n}^{2}}\delta_{\rm cn}}_{\bar{a}_{n,k}},\ \underbrace{4\pi S\mathfrak{T}\{\sigma_{tf}\}\delta_{\rm cn}}_{\bar{b}_{n,k}},\ \underbrace{8\pi^{2}S\bar{t_{n}^{2}}\delta_{\rm cn}}_{\bar{d}_{n,k}}\big{]}$ (61)
where $S$ denotes the number of samples. In addition, the remaining parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, represent geometric spreading factors, given respectively by
$[\beta_{n,k},\zeta_{n,k}]^{\rm T}=$ $\dfrac{|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}-{\bm{l}}_{n}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}-{\bm{l}}_{k}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62a)
$[\xi_{n,k},\varrho_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62b)
$[\eta_{n,k},\kappa_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\bigg{(}({\bm{l}}_{n}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{n}-{\bm{l}})}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}$
$+({\bm{l}}_{k}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{k}-{\bm{l}})}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\bigg{)}.$ (62c)
Appendix B Proof of Lemma 1
@@
It is obvious that $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ if and only if either ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$. On the other hand, in light of (42), we have
$\small\Delta L({\bm{\rho}}_{i+1})\approx\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}.$ (63)
@@
If ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ for $i>I$, we readily have $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$. Assume there exists $i>I$, where $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ with neither ${\bm{d}}_{i}\rightarrow{\bm{0}}$ nor $v_{i}\rightarrow 0$. We can solve the step size for $\Delta L({\bm{\rho}}_{i+1})=0$ as
$\small\upsilon_{i}=\dfrac{-2{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}},$ (64)
@@
Appendix C Proof of Theorem 1
@@
$\Longrightarrow\ \upsilon_{i}\leq\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}.$
where positive ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ is considered. Assuming the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded. This indicates a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus non-zero $\upsilon_{i}$. Furthermore, in light of (44), we have
$\small\Delta^{2}L({\bm{\rho}}_{i+1})\leq\epsilon_{L}^{2}\upsilon_{i}^{2}({\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}))^{2}.$ (65)
Since $L({\bm{\rho}})$ is a monotonously decreasing and bounded function regarding ${\bm{\rho}}_{i}$, the (conjugate) gradient-based descent algorithm converges to $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$. However, the deflection factor computed by (41) or the descent direction vector in (49) can guarantee a strict descent ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$. Consequently, sequence $\{L({\bm{\rho}}_{i})\}$ will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$ when it has Lipschitz gradient continuity. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when ${\bm{d}}_{i}\rightarrow{\bm{0}}$. On the other hand, we assume the objective function is not Lipschitz gradient continuous. In this case, if the sequence $\{{\bm{\rho}}_{i}\}$ pass a Lipschitz noncontinuous point ${\bm{\rho}}_{I}$ where ${\bm{d}}_{I}^{\rm T}\nabla^{2}L({\bm{\rho}}_{I}){\bm{d}}_{I}$ is extremely large, to satisfy (C), the step size should meet $\upsilon_{I}\rightarrow 0$. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ when $\upsilon_{i}\rightarrow 0$ for $i\geq I$, i.e., the sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ converge to a Lipschitz gradient noncontinuous point. $\blacksquare$